ashley99:

The equation for the area of a trapezoid is A equals one-half times h times the quantity of b subscript 1 plus b subscript 2 end quantity.. If A = 28, b1 = 6, and b2 = 8, what is the height of the trapezoid? h = 2 h = 4 h = 6 h = 7

Vocaloid:

A = (1/2)h(b1 + b2) Plug in A, b1, and b2 and solve for h

ashley99:

so it would be h=2?

Vocaloid:

28 = (1/2)(6+8)h 28 = (1/2)(14)h 28 = 7h h = 4

ashley99:

great thanks!